# A Modern Look at the Oscillation Physics Case for a Neutrino Factory

### v2

@@ 1 Introduction
The next generation of neutrino oscillation experiments, including JUNO JUNO:2022mxj , DUNE DUNE:2020ypp , and HK Hyper-KamiokandeProto-:2015xww , are expected to make significant headway in addressing three of the remaining known unknowns in particle physics. Notably, they will have excellent $\gg 5\sigma$ sensitivity to the mass ordering (MO) (the sign of $\Delta m^{2}_{31}$), they will have good $\theta_{23}$ octant determination capabilities provided that it isn’t too close to maximal, and both DUNE and HK will each individually provide $>5\sigma$ sensitivity to CP violation for significant regions of parameter space. In addition, they will provide some resolution on measuring $\delta$ (we use the standard parameterization of the PMNS Pontecorvo:1957cp ; Maki:1962mu matrix Denton:2020igp ). Atmospheric neutrino experiments such as HK, the IceCube-upgrade Ishihara:2019aao , and KM3NeT-ORCA KM3Net:2016zxf will also provide complementary information. For a recent overview of this picture, see Denton:2022een .
@@
While the outlook is hopefully rosy, it is important to carefully examine what, if anything, should come after these experiments in the neutrino oscillation program. Such an experiment would provide high precision measurements of all the oscillation parameters [^3] The solar parameters will not be well measured in long-baseline experiments Denton:2023zwa , but JUNO will provide high precision measurements of them JUNO:2022mxj and long-baseline experiments will probe the rest., test new physics scenarios in oscillations, and address any potential anomalies that arise in upcomming DUNE and HK measurements. Several experiments have been proposed with varying levels of detail. One is using a water-based liquid scintillator detector THEIA which combines the advantages of water and liquid scintillator, likely in place of one of the four LArTPC modules for DUNE Theia:2019non ; Theia:2022uyh . Another is an additional HK-like tank in Korea experiencing a different portion of the same beam that HK experiences to target the second oscillation maximum in appearance Hyper-Kamiokande:2016srs . There is also discussion about building a very large (much larger than HK) water Cherenkov detector in Sweden combined with additional beam upgrades of the ESS to create the ESSnuSB also to target the second oscillation maximum in ESSnuSB:2021azq ; Alekou:2022mav ; Abele:2022iml .
@@
In this article we discuss a different option, a neutrino factory. Neutrino factories were seriously considered as an option to probe CP violation in the first decade of this millennium largely due to the expectation of small $\theta_{13}<1^{\circ}$ which severely limits the possible size of the CP signal DeRujula:1998umv . In general, a neutrino factory is a muon storage ring in a racetrack configuration with long straight sections pointing towards the far detector and provides powerful beams of both muon and electron neutrinos of opposite charges Geer:1997iz ; Blondel:2000gj ; Blondel:2006su ; FernandezMartinez:2010zza . The initial spectrum of this decay is extremely well characterized unlike the initial spectrum from traditional fixed target neutrino oscillation experiments which primarily produce neutrinos from light meson decays with a range of energies. See Bogacz:2022xsj for a recent review of the history of neutrino factories as well as a discussion of the non-oscillation physics program at such a machine.
@@
A neutrino factory may also make sense as a first step on the path to a multi-TeV high energy muon collider to probe the electroweak sector, see e.g. Accettura:2023ked . This first step could be chronologically as a part of a technology demonstrator, a part of the accelerator chain, or both. In addition, the recent 2023 P5 report suggested consideration of neutrino physics to be done with a muon accelerator p5_2023 .
@@ 2 Neutrino Factory Setup
  • The composition and the expected energy of the neutrino beam is extremely well characterized and radiative corrections to muon decay are small Greub:1993kg ; Laing:2008zzb ; Tomalak:2021lif . Furthermore, no $\nu_{\tau}$ get produced in the source, allowing for cleaner searches for $\nu_{\tau}$ appearance channels.
  • The neutrino beam at a neutrino factory is composed of both $\nu_{\mu}$ and $\nu_{e}$ and of different charges allowing studies for channels not accessible in other controlled environments including $\nu_{e}\to\nu_{\mu}$ and $\nu_{e}\to\nu_{e}$; see Denton:2018cpu .
The first studies of a neutrino factory studies were focused on measuring the complex phase $\delta$ under the assumption of a very small $\theta_{13}$. These studies identified numerous interesting approximate degeneracies in the oscillation analysis Barger:2001yr ; Freund:2001ui ; Huber:2006wb . For small values of $\theta_{13}$, charge ID (CID) of muons or electrons was crucial to suppress the beam background for the appearance channels and to measure $\delta$. For this reason, many studies assumed a 100 kT magnetized iron detector IDS-NF:2011swj .
@@
We consider two possible neutrino factory scenarios: the neutrino source is Fermilab and, to make use of existing facilities, the far detector to be located at SURF, resulting in a baseline of 1284.9 km. Alternatively, we also consider the neutrino source at Brookhaven National Laboratory (BNL), to make use of the AGS/RHIC/EIC accelerator complex, which leads to baseline of 2542.3 km to SURF. [^4] See Kitano:2024kdv for the study of a neutrino factory based at J-PARC. We assume as matter density for both configurations 2.848 g/$\text{cm}^{3}$; the average matter density for the BNL configuration would actually be somewhat higher which would slightly enhance the differences between the two seen below. We assume that the far detector is a LArTPC configuration with a total fiducial target mass of 40 kT consistent with the envisioned DUNE detector. Based on previous exploratory studies of neutrino factories we assume $10^{21}$ muon (or anti-muon) decays per year IDS-NF:2011swj . This would require a comparable (to within $\mathcal{O}(1)$ factors) beam power as the 1.2 MW beam expected for DUNE.
@@
To simulate a neutrino factory we make use of DUNE’s far detector design DUNE:2021cuw in the GLoBES Huber:2004ka framework. We also consider oscillated tau neutrinos events with a tau neutrino CC cross section from psf2023008064 , although we treat the $\nu_{\tau}$ events as background only Donini:2010xk and don’t consider tau signal events for this three-flavor oscillation study due to the large uncertainties on the signal efficiency at LAr detectors. We have verified that including this channel would only marginally improve the sensitivity to measure the oscillation parameters because the $\nu_{\tau}$ cross section is significantly suppressed at these energies.
@@
While we do not perform a full end-to-end simulation of DUNE or a neutrino factory as that is beyond the scope of this work and would require many detector and accelerator details that are still under development, we do implement uncertainties designed to match what is generally available. We use the files provided by DUNE in DUNE:2021cuw , which largely come from DUNE:2015lol and slightly modify them. For DUNE, we take the conservative estimates for the muon and electron neutrino uncertainties of 5% and 2%, respectively. The muon neutrino uncertainty taken to be larger than the electron neutrino uncertainty because it includes determinations of the beam flux, many cross section parameters, and the fiducial volume, consistent with the approach provided by DUNE. The electron neutrino uncertainty does include a contribution due to the translation of those measurements into the appropriate values for electron neutrinos. For the beam power we take 1.2 MW with 56% uptime as recommended [^5] Note that for the neutrino factory we consider only the total number of stored muons per year, which should thus have any uptime effects already accounted for..
@@
For the backgrounds we keep with the recommendations of DUNE:2021cuw and 10% on the NC background, 20% on the tau background, and 2% on the Earth’s density.
@@ 3 Sensitivities
For our simulations we consider priors on the oscillation parameters other than $\delta$ from Gonzalez-Garcia:2021dve and we fix the MO to the normal ordering (NO), as the MO as well as all oscillation parameters will be determined via multiple independent methods with future experiments. For comparison we also show the results for true inverted ordering IO in appendix B. We focus our sensitivity studies on $\delta$ as this is the least measured parameter in neutrino oscillations, but we also investigate a neutrino factory’s sensitivity to the other relevant oscillation parameters: $\Delta m^{2}_{31}$, $\theta_{23}$, and $\theta_{13}$, while long baseline experiments only have limited sensitivity to the solar parameters Denton:2023zwa . The precision on $\delta$ does provide our primary metric quantifying the ability of a neutrino factory to differentiate among differing measurements from upcoming experiments.
Since the beam at a neutrino factory consists of equal amounts of neutrinos and anti-neutrinos, there is an inherent large contribution to the final states coming from the other charge, i.e. the processes $\nu_{e}\to\nu_{\mu}$ and $\bar{\nu}_{\mu}\to\bar{\nu}_{\mu}$ lead to the same final state although with a different energy spectrum due to different initial spectra and different oscillation probabilities, see appendix A. Differentiating the events from the disappearance channel in the analysis of the appearance channel has been identified as an important step towards a successful neutrino factory early on DeRujula:1998umv . Therefore we study the impact of CID efficiency $\epsilon_{CID}$ in terms of either electron or muon charge identification. We define the CID parameter $\epsilon_{CID}$ as
@@
with the detection efficiency $\epsilon_{f}$ and $N_{\nu}$ is the number of (anti-)neutrinos of flavor $f$ we get after taking oscillations of the initial flux into account. DUNE could potentially differentiate a muon final state from an anti-muon final state with $\epsilon_{CID}\approx 72\%$ Ternes:2019sak from muon capture on argon. In addition, inelasticity can provide some CID information as well.
Larger values of $\epsilon_{CID}$ could be achieved with magnetized detectors, an idea pursued by the INO collaboration. They estimate a CID efficiency of $\epsilon_{CID}\gtrsim 95\%$ for muons with GeV energy ICAL:2015stm . Electron CID at GeV energies has been studied in the scenario of a totally active scintillator detector in a magnetic field Bross:2007ts ; Asfandiyarov:2014haa , and a magnetized LAr detector Rubbia:1977zz ; Rubbia:2001pk ; Rubbia:2004tz ; Rubbia:2009md . Given that the final DUNE detector modules remain undesigned, a discussion of the possible benefit of working to improve CID is quite timely. In Huber:2008yx it has been pointed out that CID can also be achieved statistically at some level, provided that the energy resolution is sufficiently good.
To compare our results for a NF and to get reliable conclusions about the potential state of oscillation physics with an inclusion of a NF after the next generation of LBL experiments we combine and compare the NF results to DUNE’s and HK’s forecasted results. For DUNE we assume 480 kT-MW-year which corresponds to 5 years of each neutrino running and anti-neutrino with 1.2 MW proton beam and with a total fiducial volume of 40 kT of LAr. For HK we assume 190 kT water detector, 1.3 MW beam running for 10 years with $\nu:\bar{\nu}=1:3$ from Scott:2020gng . We simulate both experiments within the GLoBES framework.
Figure 2: The expected $1\sigma$ precision on $\delta$ as a function of the muon energy for a total of 10 years at a NF (without inputs from DUNE/HK) assuming $10^{21}$ muon decays per year, for two different baselines and two different true values of $\delta$. The colors show different assumption of CID capabilities. For comparison, the horizontal lines indicate the expected sensitivity of 10 years of DUNE.
@@ 3.1 Sensitivity to $\delta$
In fig. 2 we show the sensitivity to $\delta^{\text{true}}=(-90^{\circ},~{}0)$ as a function of muon energy for a total of $10^{22}$ muon decays and a 40 kT FD. For the FNAL-SURF baseline we find that a muon energy of $E_{\mu}\simeq 5$ GeV with equal muon and anti-muon running time allows for the most precision on the determination of $\delta$ for most cases. For the BNL-SURF baseline we find a higher energy $E_{\mu}\simeq 8$ GeV is optimal, although higher energies have similar results. The higher energy is necessary for the longer baseline to get the appearance maximum to occur within the muon decay spectrum at that baseline. In general we find that the optimal region covers a range of energies as the neutrino flux at a NF has a broad spectrum. Assuming perfect CID abilities slightly increases the optimal energy range. For these optimal energies a NF can improve over 480 kT-MW-year of DUNE assuming $10^{22}$ total muon decays. At low muon energies a NF could be sensitive to the second oscillation maximum as demonstrated by the dips in the precision curve in fig. 2, however such low energy muons have a short lifetime which could lead to technical problems in realizing this setup.
@@
To show the effect of CID in more detail we show in fig. 4 the expected $1\sigma$ precision on $\delta$ as a function of CID and exposure for the FNAL-SURF setup for $\delta_{\text{true}}=0,-90^{\circ}$. We include the anticipated information from 10 years each of DUNE and HK. Our results show that CID can improve the precision on $\delta$ by 15-20%. We have also studied the effect of assuming muon and electron CID at the same time and the results are similar. We find that electron CID has a bigger impact on the improvements in precision on $\delta$ than muon CID. This is likely due to the fact that the detected $\nu_{\mu}\to\nu_{e}$ spectrum is more similar to the detected $\bar{\nu}_{e}\to\bar{\nu}_{e}$ spectrum than the detected $\nu_{e}\to\nu_{\mu}$ spectrum to the detected $\bar{\nu}_{\mu}\to\bar{\nu}_{\mu}$ spectrum, see appendix A.
@@
We find that the sensitivity of a neutrino factory comes somewhat more from the $\nu_{e}\to\nu_{\mu}$ golden channel DeRujula:1998umv ; Cervera:2000kp and less from the $\nu_{\mu}\to\nu_{e}$ channel which will be leveraged by future proton accelerator neutrino experiments. The oscillation probabilities in these two channels are the same up to the change $\delta\to-\delta$, so each channel serves as an important cross check of the other, but provides no fundamentally new insights into the nature of CP violation or conservation. They do, however, provide a probe of CPT invariance because a neutrino factory is sensitive to four different appearance channels which CP, T, and CPT conjugates of each other. Thus a neutrino factory, especially when combined with DUNE and HK, will provide a valuable unique test of CPT invariance.
@@
Overall, the improvements on precision on $\delta$ with the BNL-SURF setup are larger than at the FNAL-SURF setup largely because in the former setup the neutrinos experience more matter effects Wolfenstein:1977ue due to the higher neutrino energy which increases the appearance probability. Furthermore, due to the higher neutrino energy in this setup the cross section uncertainty is smaller. Finally, even though the flux is lower in the BNL-SURF setup due to the larger baseline and beam spreading, the cross section is higher, the appearance probability is somewhat larger, and the variation in the probability due to variations in $\delta$ is also somewhat larger; these effects provide some cancellation for the lower flux.
@@
Figure 4: Precision on $\delta$ in degrees as a function of neutrino factory exposure and CID. The neutrino factory has a baseline and muon energy of 1284.9 km and 5 GeV and also includes 10 years of both DUNE and HK. The shaded regions are for electron CID and the lines are for muon CID; electron CID does slightly better in most cases. The horizontal lines show possibly achievable muon CID. The true values of $\delta$ are 00 (left) and $-90^{\circ}$ (right). We assume as ratio of neutrino to anti-neutrino running time of 1:1 for the neutrino factory and DUNE 1:1 and 1:3 for HK.
@@
While increased precision on $\delta$ and other oscillation parameters for the sake of precision is an important goal of particle physics, we also provide some context for these sensitivities. Because $\delta$ is the last remaining oscillation parameter that is largely unconstrained, it plays a key role in flavor model predictions Girardi:2014faa ; Everett:2019idp , along with other parameters such as $\theta_{12}$ and $\Delta m^{2}_{21}$ Gehrlein:2022nss which are expected to become measured with much more precision in the coming years by JUNO JUNO:2022mxj . While there are many conceivable predictions on neutrino parameters from many different structures (see Denton:2023hkx for a recent survey), here we focus on charged lepton corrections as they predict the mixing parameters and not the masses. In fig. 5 we show the predictions of $\cos\delta$ from four popular charged lepton correction structures as discussed in Everett:2019idp alongside the expected $1\sigma$ precision on $\cos\delta$ (taking the wider allowed ranges in all cases due to the sign degeneracy since long-baseline experiments predominantly measure $\sin\delta$ from their appearance channels, although they also probe $\cos\delta$ from their disappearance measurements Denton:2023qmd ). We consider both the precision from the combination of DUNE and HK as well as the expected improvement from including the neutrino factory at the FNAL baseline. In table 1 we document the expected precision on $\delta$ for various experiments.
@@
Figure 5: The predicted regions of $\cos\delta$ due to several popular structures within the charged-lepton correction framework, adapted from Everett:2019idp ; GR1 and GR2 are two flavor model predictions based on the golden ratio Datta:2003qg ; Everett:2008et ; Rodejohann:2008ir ; Adulpravitchai:2009bg ; Feruglio:2011qq ; Ding:2011cm , TBM is the tribimaximal flavor model prediction Wolfenstein:1978uw ; Harrison:2002er ; Harrison:2002kp ; Xing:2006xa ; He:2003rm , and HEX is based on a hexagonal structure Albright:2010ap . Along the top are the expected $1\sigma$ sensitivities to $\cos\delta$ at the various central values predicted by the models. DUNE and HK are both at 5+5 years as described in the text, as is neutrino factory with the FNAL baseline and no CID.
@@
We see that while after DUNE and HK there will be some modest model discriminating capabilities in some cases, the addition of neutrino factory will increase the capability to differentiate flavor models in some key cases. Significant additional improvement in model discrimination capability is likely to also come from JUNO’s expected improved determination of the solar parameters in coming years as well.

### v3

@@ 1 Introduction
The next generation of neutrino oscillation experiments, including JUNO [1], DUNE [2], and HK [3], are expected to make significant headway in addressing three of the remaining known unknowns in particle physics. Notably, they will have excellent $\gg 5\sigma$ sensitivity to the mass ordering (MO) (the sign of $\Delta m^{2}_{31}$), they will have good $\theta_{23}$ octant determination capabilities provided that it isn’t too close to maximal, and both DUNE and HK will each individually provide $>5\sigma$ sensitivity to CP violation for significant regions of parameter space. In addition, they will provide some resolution on measuring $\delta$ (we use the standard parameterization of the PMNS [4, 5] matrix [6]). Atmospheric neutrino experiments such as HK, the IceCube-upgrade [7], and KM3NeT-ORCA [8] will also provide complementary information. For a recent overview of this picture, see [9].
@@
While the outlook for determining the neutrino oscillation parameters is hopefully rosy, it is important to carefully examine what, if anything, should come after these experiments in the neutrino oscillation program. Such an experiment would provide high precision measurements of all the oscillation parameters [^3] The solar parameters will not be well measured in long-baseline experiments [10], but JUNO will provide high precision measurements of them [1] and long-baseline experiments will probe the rest., test new physics scenarios in oscillations, and address any potential anomalies that arise in upcoming DUNE and HK measurements. Several experiments have been proposed with varying levels of detail. One is using a water-based liquid scintillator detector THEIA which combines the advantages of water and liquid scintillator, likely in place of one of the four LArTPC modules for DUNE [11, 12]. Another is an additional HK-like tank in Korea experiencing a different portion of the same beam that HK experiences to target the second oscillation maximum in appearance [13]. There is also discussion about building a very large (much larger than HK) water Cherenkov detector in Sweden combined with additional beam upgrades of the ESS to create the ESSnuSB also to target the second oscillation maximum in [14, 15, 16].
@@
In this article we discuss a different option, a neutrino factory. Neutrino factories were seriously considered as an option to probe CP violation in the first decade of this millennium largely due to the expectation of small $\theta_{13}<1^{\circ}$ which severely limits the possible size of the CP signal [17]. In general, a neutrino factory is a muon storage ring in a racetrack configuration with long straight sections pointing towards the far detector and provides powerful beams of both muon and electron neutrinos of opposite charges [18, 19, 20, 21]. The initial spectrum of this decay is extremely well characterized unlike the initial spectrum from traditional fixed target neutrino oscillation experiments which primarily produce neutrinos from light meson decays with a range of energies. See [22] for a recent review of the history of neutrino factories as well as a discussion of the non-oscillation physics program at such a machine.
@@
A neutrino factory may also make sense as a first step on the path to a multi-TeV high energy muon collider to probe the electroweak sector, see e.g. [23]. This first step could be chronologically as a part of a technology demonstrator, a part of the accelerator chain, or both. In addition, the recent 2023 P5 report suggested consideration of neutrino physics to be done with a muon accelerator [24].
@@ 2 Neutrino Factory Setup
  • The composition and the expected energy of the neutrino beam is extremely well characterized and radiative corrections to muon decay are precisely known and small [25, 26, 27]. Furthermore, no $\nu_{\tau}$ get produced in the source, allowing for cleaner searches for $\nu_{\tau}$ appearance channels.
  • The neutrino beam at a neutrino factory is composed of both $\nu_{\mu}$ and $\nu_{e}$ and of different charges allowing studies for channels not accessible in other controlled environments including $\nu_{e}\to\nu_{\mu}$ and $\nu_{e}\to\nu_{e}$; see [28].
The first studies of a neutrino factory were focused on measuring the complex phase $\delta$ under the assumption of a very small $\theta_{13}$. These studies identified numerous interesting approximate degeneracies in the oscillation analysis [29, 30, 31]. For small values of $\theta_{13}$, charge ID (CID) of muons or electrons was crucial to suppress the beam background for the appearance channels and to measure $\delta$. For this reason, many studies assumed a 100 kT magnetized iron detector [32].
@@
We consider two possible neutrino factory scenarios: the neutrino source is Fermilab and, to make use of existing facilities, the far detector to be located at SURF, resulting in a baseline of 1284.9 km. Alternatively, we also consider the neutrino source at Brookhaven National Laboratory (BNL), to make use of the AGS/RHIC/EIC accelerator complex, which leads to baseline of 2542.3 km to SURF. [^4] See [33] for the study of a neutrino factory based at J-PARC. We assume as matter density for both configurations 2.848 g/$\text{cm}^{3}$; the average matter density for the BNL configuration would actually be somewhat higher which would slightly enhance the differences between the two seen below. We assume that the far detector is a LArTPC configuration with a total fiducial target mass of 40 kT consistent with a possible final version of the envisioned DUNE detector. Based on previous exploratory studies of neutrino factories we assume $10^{21}$ muon (or anti-muon) decays per year [32]. This would require a comparable (to within $\mathcal{O}(1)$ factors) beam power as the 1.2 MW beam expected for DUNE.
@@
To simulate a neutrino factory we make use of DUNE’s far detector design [34] in the GLoBES [35] framework. We also consider oscillated tau neutrinos events with a tau neutrino CC cross section from [36], although we treat the $\nu_{\tau}$ events as background only [37] and don’t consider tau signal events for this three-flavor oscillation study due to the large uncertainties on the signal efficiency at LAr detectors. We have verified that including this channel would only marginally improve the sensitivity to measure the oscillation parameters because the $\nu_{\tau}$ cross section is significantly suppressed at these energies.
@@
While we do not perform a full end-to-end simulation of DUNE or a neutrino factory as that is beyond the scope of this work and would require many detector and accelerator details that are still under development, we do implement uncertainties designed to match what is generally available. To simulate DUNE, we use the files provided by DUNE in [34], which largely come from [38] and slightly modify them. For DUNE, we take the conservative estimates for the muon and electron neutrino event rate normalization uncertainties of 5% and 2%, respectively. The muon neutrino uncertainty taken to be larger than the electron neutrino uncertainty because it includes determinations of the beam flux, many cross section parameters, and the fiducial volume, consistent with the approach provided by DUNE. The electron neutrino uncertainty does include a contribution due to the translation of those measurements into the appropriate values for electron neutrinos. For the beam power we take 1.2 MW with 56% uptime as recommended [^5] Note that for the neutrino factory we consider only the total number of stored muons per year, which thus have any uptime effects already accounted for.. While this run time configuration for DUNE of 480 kT-MW-years is not exactly identical to the latest DUNE staging, it is quite similar to the expected exposure after 10 years of DUNE [39, 40] and provides for an important benchmark to compare with a neutrino factory.
@@
For the backgrounds we keep with the recommendations of [34] and 10% on the normalization of the NC background, 20% on the normalization of the tau background [^6] See e.g. [41, 36] for similar numbers. We are slightly more optimistic (25$\to$20%) as this number includes a flux uncertainty measured at the near detector which should be improved at a neutrino factory., and 2% on the Earth’s density.
@@ 3 Sensitivities
For our simulations we consider priors on the oscillation parameters other than $\delta$ from [42] and we fix the MO to the normal ordering (NO), as the MO as well as all oscillation parameters will be determined via multiple independent methods with future experiments. For comparison we also show the results for true inverted ordering IO in appendix B. We focus our sensitivity studies on $\delta$ as this is the least well-measured parameter in neutrino oscillations, but we also investigate a neutrino factory’s sensitivity to the other relevant oscillation parameters: $\Delta m^{2}_{31}$, $\theta_{23}$, and $\theta_{13}$, while long baseline experiments only have limited sensitivity to the solar parameters [10]. The precision on $\delta$ does provide our primary metric quantifying the ability of a neutrino factory to differentiate among differing measurements from upcoming experiments.
Since the beam at a neutrino factory consists of equal amounts of neutrinos and anti-neutrinos, there is an inherent large contribution to the final states coming from the other charge, i.e. the processes $\nu_{e}\to\nu_{\mu}$ and $\bar{\nu}_{\mu}\to\bar{\nu}_{\mu}$ lead to the same final state although with a different energy spectrum due to different initial spectra and different oscillation probabilities, see appendix A. Differentiating the events from the disappearance channel in the analysis of the appearance channel has been identified as an important step towards a successful neutrino factory early on [17]. Therefore we study the impact of CID efficiency $\epsilon_{CID}$, which we take to be energy independent, in terms of either electron or muon charge identification. We define the CID parameter $\epsilon_{CID}$ as
@@
with the detection efficiency $\epsilon_{f}$ and $N_{\nu}$ is the number of (anti-)neutrinos of flavor $f$ we get after taking oscillations of the initial flux into account. DUNE could potentially differentiate a muon final state from an anti-muon final state with $\epsilon_{CID}\approx 72\%$ [43] from muon capture on argon. In addition, inelasticity can provide some CID information as well.
Larger values of $\epsilon_{CID}$ could be achieved with magnetized detectors, an idea pursued by the INO collaboration. They estimate a CID efficiency of $\epsilon_{CID}\gtrsim 95\%$ for muons with GeV energy [44]. Electron CID at GeV energies has been studied in the scenario of a totally active scintillator detector in a magnetic field [45, 46], and a magnetized LAr detector [47, 48, 49, 50]. Given that the final DUNE detector modules remain undesigned, a discussion of the possible benefit of working to improve CID is quite timely. In [51] it has been pointed out that CID can also be achieved statistically at some level, provided that the energy resolution is sufficiently good.
To compare our results for a NF and to get reliable conclusions about the potential state of oscillation physics with an inclusion of a NF after the next generation of LBL experiments we combine and compare the NF results to DUNE’s and HK’s forecasted results. For DUNE we assume 480 kT-MW-year which corresponds to 5 years of each neutrino running and anti-neutrino with 1.2 MW proton beam and with a total fiducial volume of 40 kT of LAr. For HK we assume 190 kT water detector, 1.3 MW beam running for 10 years with $\nu:\bar{\nu}=1:3$ from [52]. We simulate both experiments within the GLoBES framework.
Figure 2: The expected $1\sigma$ precision on $\delta$ as a function of the muon energy for a total of 10 years at a NF (without inputs from DUNE/HK) assuming $10^{21}$ muon decays per year, for two different baselines and two different true values of $\delta$ in true NO. The colors show different assumption of CID capabilities. For comparison, the horizontal lines indicate the expected sensitivity of 10 years of DUNE.
@@ 3.1 Sensitivity to $\delta$
[…]
@@
To show the effect of CID in more detail we show in fig. 4 the expected $1\sigma$ precision on $\delta$ as a function of CID and exposure for the FNAL-SURF setup for $\delta_{\text{true}}=0,-90^{\circ}$. We include the anticipated information from 10 years each of DUNE and HK. Our results show that CID can improve the precision on $\delta$ by 15-20%. We have also studied the effect of assuming muon and electron CID at the same time and the results are similar. We find that electron CID has a bigger impact on the improvements in precision on $\delta$ than muon CID. This is due to the fact that the detected $\nu_{\mu}\to\nu_{e}$ spectrum is more similar to the detected $\bar{\nu}_{e}\to\bar{\nu}_{e}$ spectrum than the detected $\nu_{e}\to\nu_{\mu}$ spectrum to the detected $\bar{\nu}_{\mu}\to\bar{\nu}_{\mu}$ spectrum, see appendix A.
@@
We find that the sensitivity of a neutrino factory comes somewhat more from the $\nu_{e}\to\nu_{\mu}$ golden channel [17, 53] and less from the $\nu_{\mu}\to\nu_{e}$ channel which will be leveraged by future proton accelerator neutrino experiments. The oscillation probabilities in these two channels are the same up to the change $\delta\to-\delta$, so each channel serves as an important cross check of the other, but provides no fundamentally new insights into the nature of CP violation or conservation. They do, however, provide a probe of CPT invariance because a neutrino factory is sensitive to four different appearance channels which CP, T, and CPT conjugates of each other. Thus a neutrino factory, especially when combined with DUNE and HK, will provide a valuable unique test of CPT invariance.
@@
Overall, the improvements on precision on $\delta$ with the BNL-SURF setup are larger than at the FNAL-SURF setup largely because in the former setup the neutrinos experience more matter effects [54] due to the higher neutrino energy which increases the appearance probability. Furthermore, due to the higher neutrino energy in this setup the cross section uncertainty is smaller. Finally, even though the flux is lower in the BNL-SURF setup due to the larger baseline and beam spreading, the cross section is higher, the appearance probability is somewhat larger, and the variation in the probability due to variations in $\delta$ is also somewhat larger; these effects provide some cancellation for the lower flux.
@@
Figure 4: Precision on $\delta$ in degrees as a function of neutrino factory exposure and CID. The neutrino factory has a baseline and muon energy of 1284.9 km and 5 GeV and also includes 10 years of both DUNE and HK. The shaded regions are for electron CID and the lines are for muon CID; electron CID does slightly better in most cases. The horizontal lines show possibly achievable muon CID. The true values of $\delta$ are 00 (left) and $-90^{\circ}$ (right). We assume as ratio of neutrino to anti-neutrino running time of 1:1 for the neutrino factory and DUNE 1:1 and 1:3 for HK. We assume true NO.
@@
While increased precision on $\delta$ and other oscillation parameters for the sake of precision is an important goal of particle physics, we also provide some context for these sensitivities. Because $\delta$ is the last remaining oscillation parameter that is largely unconstrained, it plays a key role in flavor model predictions [55, 56], along with other parameters such as $\theta_{12}$ and $\Delta m^{2}_{21}$ [57] which are expected to become measured with much more precision in the coming years by JUNO [1]. While there are many conceivable predictions on neutrino parameters from many different structures (see [58] for a recent survey), here we focus on charged lepton corrections as they predict the mixing parameters and not the masses. Charged lepton corrections start with a mass matrix for neutrinos derived from some symmetry leading to popular and well-known structures such as tri-bimaximal (TBM) which predicts $\theta_{12}\approx 35^{\circ}$, $\theta_{23}=45^{\circ}$, and $\theta_{13}=0$, and thus no CP violation. A small correction is thus needed to explain $\theta_{13}=8.5^{\circ}$ which also allows for CP violation, a slightly smaller value of $\theta_{12}$, and possible variation in $\theta_{23}$; these are usually provided in the charged lepton mass matrix which, in addition to the neutrino mass matrix, provides a slightly off-diagonal matrix parameterized by a Cabibbo sized angle. In fig. 5 we show the predictions of $\cos\delta$ from four popular charged lepton correction structures as discussed in [56] alongside the expected $1\sigma$ precision on $\cos\delta$ (taking the wider allowed ranges in all cases due to the sign degeneracy since long-baseline experiments predominantly measure $\sin\delta$ from their appearance channels, although they also probe $\cos\delta$ from their disappearance measurements [59]). We consider both the precision from the combination of DUNE and HK as well as the expected improvement from including the neutrino factory at the FNAL baseline. In table 1 we document the expected precision on $\delta$ for various experiments.
@@
Figure 5: The predicted regions of $\cos\delta$ due to several popular structures within the charged-lepton correction framework, adapted from [56]; GR1 and GR2 are two flavor model predictions based on the golden ratio [60, 61, 62, 63, 64, 65], TBM is the tribimaximal flavor model prediction [66, 67, 68, 69, 70], and HEX is based on a hexagonal structure [71]. Along the top are the expected $1\sigma$ sensitivities to $\cos\delta$ at the various central values predicted by the models. DUNE and HK are both at 5+5 years as described in the text, as is neutrino factory with the FNAL baseline and no CID.
@@
We see that while after DUNE and HK there will be some modest model discriminating capabilities in some cases, the addition of neutrino factory will somewhat increase the capability to differentiate flavor models in some standard benchmark cases as shown in fig. 5. Significant additional improvement in model discrimination capability is likely to also come from JUNO’s expected improved determination of the solar parameters in coming years as well.